$SPAD/src/input f07aef.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
trans:="N"
n:=4
nrhs:=2
a:Matrix SF:=
 [[5.25 ,-2.95 ,-0.95 ,-3.8 ],_
  [0.34 ,3.89 ,2.38 ,0.41 ],_
  [0.3 ,-0.46 ,-1.51 ,0.29 ],_
  [-0.21 ,-0.33 ,0.01 ,1.13 ]]
lda:=4
ipiv:Matrix Integer:=
 [[2 ,2 ,3 ,4 ]]
ldb:=4
b:Matrix SF:=
  [[9.52 ,18.47 ],_
   [24.35 ,2.25 ],_
   [0.77 ,-13.3 ],_
   [-6.22 ,-6.21 ]]
result:=f07aef(trans,n,nrhs,a,lda,ipiv,ldb,b)

———-

References^{[1] nothing}

4